### v1

@@ 4.3 Power allocation strategies
In order to keep the system complexity at a reasonable level and to come up with a system implementation practically deployable, a simple fractional power allocation (FPA) strategy is considered here. We denote as $P_{m,{\rm max}}^{(a)}$ and $P_{\ell,{\rm max}}^{(b)}$ the maximum powers available at the $m$-th AP and at the $\ell$-th BS for the downlink transmission, respectively. The generic $m$-th AP ($\ell$-th BS) serves the UEs in $\mathcal{K}_{m}$ ($\mathcal{J}_{\ell}$), i.e. those UEs such that $a_{k,m}=1$ ($b_{k,\ell}=1$), with the following power
@@
In (15), $\alpha$ is a real parameter, usually taking value in the range $[-1,1]$. Stategies (15) define thus a family of power control strategies, depending on the chosen value of $\alpha$. Notice that for $\alpha=0$ we have the simple uniform power allocation, where each AP/BS equally divides the available power among the UEs that he is serving. Positive values of $\alpha$ usually privilege fairness across users (i.e., more power is given to UEs with weaker channels), whereas negative values of $\alpha$ are beneficial for the UEs with the best channel conditions.
@@ 5 FULL-Coop scenario: joint precoding scheme
We now consider, for the FULL-Coop scenario a joint precoding scheme exploiting the cooperation between BSs and APs. Indeed, in practical implementations, the centralized computation of the beamformer must take into account the local power budget constraints at each AP and each BS. In the following, we provide a procedure so that this constraint can be fulfilled. For the sake of brevity, we now discuss only the joint partial zero-forcing (JPZF) but all the precoding schemes discussed above for the local processing can be considered. For computing the centralized beamformer, one scaling factor must be used for each UE in the network. We thus assume that:
@@
To effectively determine $\mathbf{w}_{k}^{\rm JPZF}$, we first form the $\left[\left(N_{\rm AP}\widetilde{N}_{\rm AP}+N_{\rm BS}\widetilde{N}_{\rm BS}\right)\times R_{\rm JPZF}\right]$-dimensional matrix $\mathbf{I}_{k}^{\rm JPZF}$, whose column are the vectors $\left\{\widetilde{\mathbf{u}}_{j}^{(k)}\right\}_{j\in\mathcal{R}_{k}}$. Then, the $\left(N_{\rm AP}\widetilde{N}_{\rm AP}+N_{\rm BS}\widetilde{N}_{\rm BS}\right)$-dimensional beamforming vector $\mathbf{w}_{k}^{\rm JPZF}$ can be obtained as
@@
In order to allocate the same power to all the users and ensure that the maximum transmit power per AP and per BS is fulfilled, the following procedure is adopted. We collect the beamforming vectors in the $\left[\left(N_{\rm AP}M+N_{\rm BS}L\right)\times K\right]$-dimensional matrix $\mathbf{Q}$ using the pseudo code reported in Algorithm 2. As we can note from Algorithm 2, matrix $\mathbf{Q}$ contains the partition of the joint precoding vector transmitted by each AP and BS in the cooperation network.
@@
Upon defining the matrix $\mathbf{Q}$, the equal stream power allocation can be solved as follows. Upon defining the $K$-dimensional vector $\bm{\eta}=\left[\eta_{1},\ldots,\eta_{K}\right]^{T}$, the following power constraints must be satisfied
@@ 6 FULL-Coop scenario: consideration of limited-capacity fronthaul constraint
Considering again the FULL-Coop scenario, we now provide an heuristic procedure aimed at the reduction of the fronthaul load. We consider both the local processing at the AP/BS discussed in Section 4.2 and the joint processing presented in Section 5, and we take into account also the use of multiple subcarriers. We assume that the generic $n$-th AP/BS is connected to the CPU via a fronthaul link with a capacity of $\bar{F}_{n},\forall n$. First of all, we formulate the total FH requirement at the $n$-th AP/BS as the sum of the FH data and FH weights requirements, and we aim at designing a UE association algorithm such that the fronthaul link capacity is not exceeded, i.e.,
@@
Under the assumption that all the users are served over all the subcarriers, if the $k$-th UE is served by the $n$-th AP/BS, the CPU needs to send over the FH link to the $n$-th AP/BS the $k$-th UE information symbols for all the subcarriers. Accordingly, the data rate related to the FH data is written as
@@
where $M_{\rm QAM}$ is the modulation cardinality, $N_{\rm RB}$ is the number of resource blocks, $N_{\rm sc}^{(RB)}$ is the number of subcarriers per resource block, $N_{\rm sc}^{(OFDM)}$ is the number of OFDM symbols per resource block, $\tau_{\rm data}$ is the transmit delay for the data, and $\eta_{\rm CPRI}$ is the efficiency of the Common Public Radio Interface (CPRI) [^4] Numerical values used for these parameters in the numerical results will be reported later in Section 7.1.. The data-rate related to the FH weigths depends on the number of bits needed to send the beamforming weigths to the $n$-th AP/BS. Note that this value is zero in the case of local processing since in this case the beamforming weights are evaluated locally at the AP/BS and thus they do not need to be transferred over the fronthaul link. We have:
@@
    Initialize the procedure by creating a, association of APs and BSs to the users assuming that each UE is served by the best $\widetilde{N}_{\rm AP}$ APs and $\widetilde{N}_{\rm AP}$ BSs. This association is encoded in the coefficients $\widetilde{a}_{k,n}\forall k,n$, with $\widetilde{a}_{k,n}=1$ if the $k$-th user is served by the $n$-th AP/BS and 0 otherwise.
@@
    Compute the precoders. Specifically, in the case of local processing discussed in Section 4.2 design the local beamforming at the APs/BSs; in the case of joint processing presented in Section 5, design the coordinate beamforming at the CPU to be shared over the fronthaul link. This procedure is done every $N_{C,B}^{(RB)}$ subcarriers using the averaged channels over the consecutive $N_{C,B}^{(RB)}$ subcarriers.
@@
    $S_{k,\overline{n}}=\displaystyle\frac{\displaystyle\sum_{\begin{subarray}{c}n=1\\ n\neq\overline{n}\end{subarray}}^{M+L}\widetilde{a}_{k,n}\delta_{k,n}}{\displaystyle\sum_{\begin{subarray}{c}j=1\\ j\neq k\end{subarray}}^{K}\displaystyle\sum_{\begin{subarray}{c}n=1\\ n\neq\overline{n}\end{subarray}}^{M+L}\widetilde{a}_{j,n}\delta_{k,n}+\sigma^{2}_{z}}\;,$
    where $\delta_{k,n}$ is the large scale fading coefficient between the $k$-th user and the $n$-th AP/BS. Notice that the metric $S_{k,\overline{n}}$ is a proxy (i.e., an approximation) of the $k$-th UE SINR for the case in which is is no longer served by the $n$-th AP/BS.
@@
    $k^{*}=\arg\max_{k\in\mathcal{K}_{\overline{n}}}S_{k,\overline{n}}$
@@ 7.1 Simulation setup and parameters
A cluster of three cells surrounded by nine cells to take into account the effect of the inter-cell interference is considered. Each cell is actually made of three sectors with $22\pi/$ 2 coverage; each of these sectors is treated as an independent base station and is equipped with a uniform linear antenna array with half wavelength spacing. The positions of the users are uniformly generated in each cell in the annulus with minimal radius (i.e. distance from a BS) 15m and maximal radius set at $0.97\text{ISD}/2$, where ISD is the ”inter-site distance,” i.e. the distance between BSs. We consider two distributions for the APs: uniformly distributed and cell-edge placement. In the former case the APs are randomly and uniformly distributed similarly to the users, while in the latter scenario the APs are placed on a circle with radius $0.8\,\text{ISD}/2$. In Fig. 1 we report a snapshot of the 2D view of the considered scenario with $\text{ISD}=500$m.
With regard to the propagation channels, we use the model in reference [32]. We start by describing the channel between the BSs and the users, which are modeled according to the urban macro cell (UMa) non-line-of-sight (NLOS) model, using the parameters reported in [32, Table 7.4.1-1].
@@
In particular, recalling Eq. (1), the large scale fading coefficient between the $k$-th user and the $\ell$-th BS in is $\rho_{k,\ell}=u_{k,\ell}\widetilde{\rho}_{k,\ell}$. Here, $u_{k,\ell}$ takes into account the array gain of the ULA at the BS. We assume that the array gain is modeled with a Gaussian shaping radiating onto a $[-\pi/2,\pi/2]$ angular region as reported in Fig. 2. The large scale fading coefficient $\widetilde{\rho}_{k,\ell}$ in logarithmic unit is evaluated as follows
@@
where $\text{PL}_{k,\ell}^{(b)}$ represents the path loss and $z_{k,\ell}^{(b)}$ represents the shadow fading. Details on the path loss modeling are reported in the appendix. The shadow fading coefficients from a BS to different users are correlated and follow [32, Eq. (7.4-5)]
@@
where $\bar{d}_{k,j}^{(b)}$ is the distance between the $k$-th and the $j$-th outdoor users and $d_{\rm corr}^{(b)}=50$m is the correlation distance and follows [32, Table 7.5-6 Part-1] and $\sigma_{(b),{\rm SF}}=6$ [32, Table 7.4.1-1].
For the channels between the APs and the users we use the urban micro cell (UMi) non-line-of-sight (NLOS) model. In particular, the path-loss between the $k$-th user and the $m$-th AP is modeled according to [32, Table 7.4.1-1]. The large scale fading coefficient $\beta_{k,m}$ in Eq. (1) is evaluated in logarithmic unit as follows
@@
where $\text{PL}_{k,m}^{(a)}$ is the path loss (described in the appendix) and $z_{k,m}^{(a)}$ represents the shadow fading. The shadow fading coefficients from an AP to different users are correlated with an expression similar to (28), where the values $\bar{d}_{k,j}^{(b)}$ equals now 13m $\sigma_{(b),{\rm SF}}=7.82$.
The chosen carrier frequency is $f_{c}=3.5$ GHz, the system bandwidth is $B=20$ MHz, the power spectral density of the noise is $N_{0}=-174$ dBm/Hz and the noise figure at the receiver is $F=9$ dB. We consider in the whole simulation area $Q=12$ macro BSs with three sectors, i.e., the total number of BSs is $L=36$, the number of APs is $M=108$, 9 in each cell. We equip the BSs ULA with 32 antennas and the APs with 8 antennas arrays; the height of the BSs is $h_{\rm BS}=25$m and the height of the APs is $h_{\rm AP}=10$m. We assume $\tau_{c}=640$ time/frequency samples, corresponding to a coherence bandwidth of $640$ kHz and a coherence time of $1$ ms. Equal uplink/downlink split of the available time/frequency resources after training is assumed, i.e., $\tau_{\rm d}=\frac{\tau_{c}-\tau_{p}}{2}$. We take a set $\mathcal{P}_{\tau_{p}}$ of orthogonal pilots with length $\tau_{p}=32$ and the pilot sequences in $\mathcal{P}_{\tau_{p}}$ are assigned to the users according to the $k$-means Algorithm discussed in Section 2.4, i.e., the results account for the impact of pilot contamination. The uplink transmit power during training is $\eta_{k}=\tau_{p}\overline{\eta}_{k}$, with $\overline{\eta}_{k}=300$ mW $\forall k=1,\ldots,K$. We present results for the case of LMMSE channel estimation discussed in Section 2.3. All the system parameters are summarized in Table 1.
@@
AP distribution Distribution of APs in the system Uniform or Cell-Edge placement
MS distribution Distribution of MSs in the system Uniform placement
@@
Figure 3: CDF of the rate per user in the four considered scenarios with FPA with $\alpha=-0.5$, LMMSE beamfoming at the BSs and APs and 5 users per sector, i.e., $K=180$. Continuous lines represent uniform APs placement in the cells and dashed lines represent cell-edge APs placement.

### v2

@@ 4.3 Power allocation strategies
In order to keep the system complexity at a reasonable level and to come up with a practically deployable system implementation, a simple fractional power allocation (FPA) strategy is considered here. We denote as $P_{m,{\rm max}}^{(a)}$ and $P_{\ell,{\rm max}}^{(b)}$ the maximum powers available at the $m$-th AP and at the $\ell$-th BS for the downlink transmission, respectively. The generic $m$-th AP ($\ell$-th BS) serves the UEs in $\mathcal{K}_{m}$ ($\mathcal{J}_{\ell}$), i.e. those UEs such that $a_{k,m}=1$ ($b_{k,\ell}=1$), with the following power
@@
In (18), $\alpha$ is a real parameter, which usually takes a value in the range $[-1,1]$. Stategies (18) define thus a family of power control strategies, depending on the chosen value of $\alpha$. Notice that for $\alpha=0$ we have the simple uniform power allocation, where each AP/BS equally divides the available power among the UEs he is serving. Positive values of $\alpha$ usually favor fairness between users (that is, more power is given to UEs with weaker channels), while negative values of $\alpha$ are beneficial for UEs with the best channel conditions.
@@ 5 FULL-Coop scenario: joint precoding scheme
We now consider, for the FULL-Coop scenario, a joint precoding scheme exploiting the cooperation between BSs and APs. In fact, in practical implementations, the centralized computation of the beamformer must take into account the local power budget constraints at each AP and each BS. In the following, we provide a procedure for fulfilling this constraint. For the sake of brevity, we now discuss only the joint partial zero-forcing (JPZF) but all the precoding schemes discussed above for the local processing can be considered. For computing the centralized beamformer, one scaling factor must be used for each UE in the network. We thus assume that:
@@
To effectively determine $\mathbf{w}_{k}^{\rm JPZF}$, we first form the matrix of dimension $\left[\left(N_{\rm AP}\widetilde{N}_{\rm AP}+N_{\rm BS}\widetilde{N}_{\rm BS}\right)\times R_{\rm JPZF}\right]$ $\mathbf{I}_{k}^{\rm JPZF}$, whose columns are the vectors $\left\{\widetilde{\mathbf{u}}_{j}^{(k)}\right\}_{j\in\mathcal{R}_{k}}$. Then, the $\left(N_{\rm AP}\widetilde{N}_{\rm AP}+N_{\rm BS}\widetilde{N}_{\rm BS}\right)$-dimensional beamforming vector $\mathbf{w}_{k}^{\rm JPZF}$ can be obtained as
@@
In order to allocate the same power to all the users and ensure that the maximum transmit power per AP and per BS is fulfilled, the following procedure is adopted. We collect the beamforming vectors in the $\left[\left(N_{\rm AP}M+N_{\rm BS}L\right)\times K\right]$-dimensional matrix $\mathbf{Q}$ using the pseudo-code reported in Algorithm 2. As we can note from Algorithm 2, matrix $\mathbf{Q}$ contains the partition of the joint precoding vector transmitted by each AP and BS in the cooperation network.
@@
Upon defining the matrix $\mathbf{Q}$, the equal-stream power allocation can be solved as follows. Upon defining the $K$-dimensional vector $\bm{\eta}=\left[\eta_{1},\ldots,\eta_{K}\right]^{T}$, the following power constraints must be satisfied
@@
5.1 Computational complexity for pre-coder computation
In the following, we evaluate the complexity of the precoder computation in terms of complex multiplications required. The calculation of $\mathbf{w}_{k}^{{\rm JPZF}}$ requires the calculation of $\widetilde{\mathbf{I}}_{k}^{\rm JPZF}=\text{orth}\left[\mathbf{I}_{k}^{\rm JPZF}\right]$, i.e., the matrix containing on its columns an orthonormal basis for the column span of the $\left[\left(N_{\rm AP}\widetilde{N}_{\rm AP}+N_{\rm BS}\widetilde{N}_{\rm BS}\right)\times R_{\rm JPZF}\right]$-dimensional matrix $\mathbf{I}_{k}^{\rm JPZF}$. This operation can be shown to require a number of complex multiplications equal to $O\left[R_{\rm JPZF}^{2}\left(N_{\rm AP}\widetilde{N}_{\rm AP}+N_{\rm BS}\widetilde{N}_{\rm BS}\right)\right]$. Then, a projection of an $\left(N_{\rm AP}\widetilde{N}_{\rm AP}+N_{\rm BS}\widetilde{N}_{\rm BS}\right)$ dimensional vector onto the orthogonal complement to the subspace spanned by the columns of $\widetilde{\mathbf{I}}_{k}^{\rm JPZF}$ is to be performed. This operation requires $O\left[R_{\rm JPZF}\left(N_{\rm AP}\widetilde{N}_{\rm AP}+N_{\rm BS}\widetilde{N}_{\rm BS}\right)\right]$ complex multiplications. The dominant term determining the complexity for the computation of the centralized precoders is thus $O\left[R_{\rm JPZF}^{2}\left(N_{\rm AP}\widetilde{N}_{\rm AP}+N_{\rm BS}\widetilde{N}_{\rm BS}\right)\right]$.
In contrast, in the case of local processing, the precoders $\mathbf{w}_{k,\ell}^{(b),{\rm PZF}}$ are calculated in $\widetilde{N}_{\rm BS}$ BS and the precoders $\mathbf{w}_{k,m}^{(a),{\rm PZF}}$ are calculated in $\widetilde{N}_{\rm AP}$ APs, leading to the cost of complex multiplications $\widetilde{N}_{\rm BS}O\left(\left[N_{\rm PZF}^{\rm BS}\right]^{2}N_{\rm BS}\right)+\widetilde{N}_{\rm AP}O\left(\left[N_{\rm PZF}^{\rm AP}\right]^{2}N_{\rm AP}^{2}\right)$.
As an example, assuming $N_{\rm BS}=32$, $N_{\rm AP}=8$, $N_{\rm PZF}^{\rm BS}=N_{\rm BS}/2=16$, $N_{\rm PZF}^{\rm AP}=N_{\rm AP}/2=4$, $\widetilde{N}_{\rm BS}=3$, $\widetilde{N}_{\rm AP}=6$, $R_{\rm JPZF}=\left(N_{\rm AP}\widetilde{N}_{\rm AP}+N_{\rm BS}\widetilde{N}_{\rm BS}\right)/2=72$, the computational complexity of the centralized precoding is $O\left[72^{2}\cdot 144\right]\approx O((7.4\cdot 10^{5}))$ complex multiplications while the complexity of the local precoding is $3O\left(16^{2}\cdot 32\right)+6O\left(4^{2}\cdot 8\right)\approx\;O(2.5\cdot 10^{4})$ complex multiplications.
@@ 6 FULL-Coop scenario: consideration of limited-capacity fronthaul constraint
Considering again the FULL-Coop scenario, we now provide a heuristic procedure aimed at reducing the fronthaul load. We consider both the local processing at the AP/BS discussed in Section 4.4.2 and the joint processing presented in Section 5, and we also take into account the use of multiple subcarriers. We assume that the generic $n$-th AP/BS is connected to the CPU via a fronthaul link with a capacity of $\bar{F}_{n},\forall n$. First of all, we formulate the total FH requirement at the $n$-th AP/BS as the sum of the FH data and FH weights requirements, and we aim at designing a UE association algorithm such that the fronthaul link capacity is not exceeded, i.e.,
@@
Under the assumption that all users are served over all the subcarriers, if the $k$-th UE is served by the $n$-th AP/BS, the CPU needs to send over the FH link to the $n$-th AP/BS the $k$-th UE information symbols for all the subcarriers. Accordingly, the data rate related to the FH data is written as
@@
where $M_{\rm QAM}$ is the modulation cardinality, $N_{\rm RB}$ is the number of resource blocks, $N_{\rm sc}^{(RB)}$ is the number of subcarriers per resource block, $N_{\rm sc}^{(OFDM)}$ is the number of OFDM symbols per resource block, $\tau_{\rm data}$ is the transmit delay for the data, and $\eta_{\rm CPRI}$ is the efficiency of the Common Public Radio Interface (CPRI) [^5] Numerical values used for these parameters in the numerical results will be reported later in Section 7.7.1.. The data-rate related to the FH weigths depends on the number of bits needed to send the beamforming weigths to the $n$-th AP/BS. Note that this value is zero in the case of local processing since in this case the beamforming weights are evaluated locally at the AP/BS and thus they do not need to be transferred over the fronthaul link. We have:
@@
    Initialize the procedure by creating an association of APs and BSs to the users assuming that each UE is served by the best $\widetilde{N}_{\rm AP}$ APs and $\widetilde{N}_{\rm AP}$ BSs. This association is encoded in the coefficients $\widetilde{a}_{k,n},\;\forall k,n$, with $\widetilde{a}_{k,n}=1$ if the $k$-th user is served by the $n$-th AP/BS and 0 otherwise.
@@
    Compute the precoders. Specifically, in the case of local processing discussed in Section 4.4.2 design the local beamforming at the APs/BSs; in the case of joint processing presented in Section 5, design the coordinate beamforming at the CPU to be shared over the fronthaul link. This procedure is performed on every $N_{C,B}^{(RB)}$ subcarrier using the averaged channels on consecutive $N_{C,B}^{(RB)}$ subcarriers.
@@
    $S_{k,\overline{n}}=\displaystyle\frac{\displaystyle\sum_{\begin{subarray}{c}n=1\\ n\neq\overline{n}\end{subarray}}^{M+L}\widetilde{a}_{k,n}\delta_{k,n}}{\displaystyle\sum_{\begin{subarray}{c}j=1\\ j\neq k\end{subarray}}^{K}\displaystyle\sum_{\begin{subarray}{c}n=1\\ n\neq\overline{n}\end{subarray}}^{M+L}\widetilde{a}_{j,n}\delta_{k,n}+\sigma^{2}_{z}}\;,$ (30)
    where $\delta_{k,n}$ is the large-scale fading coefficient between the $k$-th user and the $n$-th AP/BS. Notice that the metric $S_{k,\overline{n}}$ is a proxy (i.e., an approximation) of the $k$-th UE SINR for the case in which is is no longer served by the $n$-th AP/BS.
@@
    $k^{*}=\arg\max_{k\in\mathcal{K}_{\overline{n}}}S_{k,\overline{n}}$ (31)
@@
Figure 2: Flowchart of the precoding design and user association considering limited-capacity fronthaul constraints.
Figure 2 presents the detailed flow chart for the procedure mentioned earlier and described in Algorithm 3.
@@ 7.1 Simulation setup and parameters
A cluster of three cells surrounded by nine cells is considered to take into account the effect of intercell interference. Each cell is actually made of three sectors with $22\pi/$ 2 coverage; each of these sectors is treated as an independent base station and is equipped with a ULA with half-wavelength spacing. The positions of the users are uniformly generated in each cell in the annulus with minimal radius (that is, distance from a BS) 15 m and maximal radius set at $0.97\text{ISD}/2$, where ISD is the “inter-site distance”, i.e. the distance between BSs. We consider two distributions for the APs: uniformly distributed and cell-edge placement. In the former case, the APs are randomly and uniformly distributed similarly to the users, while in the latter scenario the APs are placed on a circle with radius $0.8\,\text{ISD}/2$. In Fig. 1 we report a snapshot of the 2D view of the scenario considered with $\text{ISD}=500$ m.
With regard to the model for the large-scale fading coefficients, we use the model in reference [35]. We start by describing the channel between the BSs and the users, which are modeled according to the urban macrocell (UMa) model, using the parameters reported in [35, Table 7.4.1-1]. In order to discriminate between LOS and non-line-of-sight (NLOS) model, we firstly evaluate the LOS probability according to the parameters reported in in [35, Table 7.4.2-1]. For the UMa scenario, the LOS probability depends on the 2D distance between the $k$-th user and the $\ell$-th BS, $d_{k,\ell}^{(b)}$, according to the expression in Eq. (33) at the top of next page, where $h_{k}$ is the antenna height at the $k$-th user, and
$C\left(h_{k}\right)=\left\{\begin{array}[]{cc}0&\quad\text{if}\;h_{k}\leq 13\,\text{m}\\ \left(\displaystyle\frac{h_{k}-13}{10}\right)^{1.5}&\quad\text{if}\;13<h_{k}\leq 23\,\text{m}\end{array}\right..$ (32)
$p_{\rm LOS}^{(b)}\left(d_{k,\ell}^{(b)}\right)=\left\{\begin{array}[]{cc}1&\quad\text{if}\;d_{k,\ell}^{(b)}\leq 18\,\text{m}\\ \left[\displaystyle\frac{18}{d_{k,\ell}^{(b)}}+\left(1-\displaystyle\frac{18}{d_{k,\ell}^{(b)}}\right)e^{-\frac{d_{k,\ell}^{(b)}}{63}}\right]\left(1+\displaystyle\frac{5}{4}C\left(h_{k}\right)\left(\frac{d_{k,\ell}^{(b)}}{100}\right)^{3}e^{-\frac{d_{k,\ell}^{(b)}}{150}}\right)&\quad\text{otherwise}\end{array}\right.$ (33)
@@
Then, recalling Eq. (2), the large-scale fading coefficient between the $k$-th user and the $\ell$-th BS in is $\rho_{k,\ell}=u_{k,\ell}\widetilde{\rho}_{k,\ell}$. Here, $u_{k,\ell}$ takes into account the gain of the ULA array in the BS. We assume that the array gain is modeled with a Gaussian shaping radiating onto a $[-\pi/2,\pi/2]$ angular region as reported in Fig. 3. The large-scale fading coefficient $\widetilde{\rho}_{k,\ell}$ in logarithmic unit is evaluated as follows
@@
where $\text{PL}_{k,\ell}^{(b)}$ represents the path loss and $z_{k,\ell}^{(b)}$ represents the shadow fading. Details on the path loss modeling are reported in the Appendix. The shadow fading coefficients from a BS to different users are correlated and follow [35, Eq. (7.4-5)]
@@
where $\bar{d}_{k,j}^{(b)}$ is the distance between the $k$-th and the $j$-th outdoor users and $d_{\rm corr}^{(b)}=50$m is the correlation distance and follows [35, Table 7.5-6 Part-1] and $\sigma_{(b),{\rm SF}}=6$ [35, Table 7.4.1-1].
For the channels between the APs and the users, we use the urban microcell (UMi) model. Again,in order to discriminate between LOS and NLOS model, we first evaluate the probability of LOS according to the parameters reported in [35, Table 7.4.2-1]. For the UMi scenario, the LOS probability depends on the 2D distance between the $k$-th user and the $m$-th AP, $d_{k,m}^{(a)}$, according to the expression in Eq. (36) at the top of next page.
$p_{\rm LOS}^{(a)}\left(d_{k,m}^{(a)}\right)=\left\{\begin{array}[]{cc}1&\quad\text{if}\;d_{k,m}^{(a)}\leq 18\,\text{m}\\ \displaystyle\frac{18}{d_{k,m}^{(a)}}+\left(1-\displaystyle\frac{18}{d_{k,m}^{(a)}}\right)e^{-\frac{d_{k,m}^{(a)}}{36}}&\quad\text{otherwise}\end{array}\right.$ (36)
Then, the path-loss between the $k$-th user and the $m$-th AP is modeled according to [35, Table 7.4.1-1]. The large-scale fading coefficient $\beta_{k,m}$ in Eq. (1) is evaluated in logarithmic unit as follows
@@
where $\text{PL}_{k,m}^{(a)}$ is the path loss (described in the Appendix) and $z_{k,m}^{(a)}$ represents the shadow fading. The shadow fading coefficients from an AP to different users are correlated with an expression similar to (35), where the values $\bar{d}_{k,j}^{(b)}$ now equal 13m, and $\sigma_{(b),{\rm SF}}=7.82$.
The chosen carrier frequency is $f_{c}=3.5$ GHz, the system bandwidth is $B=20$ MHz, the power spectral density of the noise is $N_{0}=-174$ dBm/Hz and the noise figure at the receiver is $F=9$ dB. We consider in the whole simulation area $Q=12$ macro BSs with three sectors, i.e., the total number of BSs is $L=36$, the number of APs is $M=108$, 9 in each cell. We equip the BSs ULA with 32 antennas and the APs with 8 antennas arrays; the height of the BSs is $h_{\rm BS}=25$ m and the height of the APs is $h_{\rm AP}=10$ m. We assume $\tau_{c}=640$ time/frequency samples, corresponding to a coherence bandwidth of $640$ kHz and a coherence time of $1$ ms. Equal uplink/downlink split of the available time/frequency resources after training is assumed, i.e., $\tau_{\rm d}=\frac{\tau_{c}-\tau_{p}}{2}$. We take a set $\mathcal{P}_{\tau_{p}}$ of orthogonal pilots with length $\tau_{p}=32$ and the pilot sequences in $\mathcal{P}_{\tau_{p}}$ are assigned to the users according to the $k$-means Algorithm discussed in Section 2.2.4, i.e., the results account for the impact of pilot contamination. The uplink transmit power during training is $\eta_{k}=\tau_{p}\overline{\eta}_{k}$, with $\overline{\eta}_{k}=300$ mW $\forall k=1,\ldots,K$. We present results for the case of LMMSE channel estimation discussed in Section 2.2.3. All the system parameters are summarized in Table 2.
@@
AP dist. Distribution of APs in the system Uniform or Cell-Edge placement
MS dist. Distribution of MSs in the system Uniform placement
@@
$\widetilde{N}_{\rm BS}$ number of BSs serving each user in the FULL-Coop scenario 3
$\widetilde{N}_{\rm AP}$ number of APs serving each user in the FULL-Coop scenario 6
$N_{\rm PZF}^{\rm BS}$ number of users to be protected by the interference generated by each BS $N_{\rm BS}/2=16$
$N_{\rm PZF}^{\rm AP}$ number of users to be protected by the interference generated by each AP $N_{\rm AP}/2=4$
$R_{\rm JPZF}$ number of users to be protected from the interference generated by the downlink transmission to the generic user in the centralized precoding scheme $\begin{array}[]{lll}\!\!\!\!\!\!\left(N_{\rm AP}\widetilde{N}_{\rm AP}+\right.\\ \!\!\!\!\!\!\!\!\!\left.N_{\rm BS}\!\widetilde{N}_{\rm BS}\right)\!\!/2\!=\!72\end{array}$
@@
Figure 4: CDF of the rate per user in the four considered scenarios with FPA with $\alpha=-0.5$, MMSE beamfoming at the BSs and APs and 5 users per sector, i.e., $K=180$. Continuous lines represent uniform APs placement in the cells and dashed lines represent cell-edge APs placement.